Tim discovered that 2 out of every 45 toy trucks made at the county toy factory is broken. On a large order of 180 trucks, how many broken trucks can Tim likely expect to find?
Answer by stanbon(75887)

Tim discovered that 2 out of every 45 toy trucks made at the county toy factory is broken. On a large order of 180 trucks, how many broken trucks can Tim likely expect to find?
------
x/180 = 2/45
----
x = 180(2/45)
----
x = 8
===============
Cheers,
stan H.
==============